# Localized Multi-Dimensional Patterns

### v1

@@ 1 Introduction
In this review, we will focus on structures that are localized in one or more spatial dimensions embedded in a quiescent state and possess some non-trivial fine-scale regular structure, which we call a pattern. Figure 1 provides examples of the types of localized patterns we will concentrate on that occur in both continuum and discrete systems. Throughout this review, we will see that the field has been developed with numerical, experimental, and analytical techniques that work in tandem to better understand localization mechanisms. We highlight the three reviews that came before this one [74, 138, 137], but note their focus is more rooted in the physics literature or primarily restricted to localized patterns in one spatial dimension. What differentiates this review from those that came before is that we seek to highlight the mathematical development of the subject. This includes how specific equations have been investigated for localized patterns and providing overviews of the myriad of clever mathematical and computational techniques that have been developed to study these multi-dimensional patterns.
Localized patterns can be found in a wide range of application areas from fluids, quantum theory, biology, mechanics, sociology, to nonlinear optics. They come in the form of an oasis of vegetation in an arid landscape, a hotspot of criminal activity, or a laser pulse in plasmas. Because they arise in so many different applications, they often are referred to by different names which have historical relevance to the field. This includes solitary waves in fluid dynamics, convectons in binary fluid convection, and fairy circles in vegetation models, to name a few. To ease the reader’s journey into the literature on localized patterns, we will provide references and nomenclature for each specific type of pattern as they occur in this review. This will help to connect the specific pattern to specific applications. However, we note that while the study of these patterns is motivated by application, our goal is to review all things mathematical in their investigation. This means primarily relying on phenomenological equations that describe macroscopic behavior and attempting to understand the fundamental mechanisms that drive their formation.
@@
where $\mathbf{u}=\mathbf{u}(\mathbf{x},t)\in\mathbb{R}^{n}$, $\Delta$ is the Laplace operator on $\mathbb{R}^{d}$, $\mathbf{f}$ is a smooth nonlinearity, $\mu$ is a bifurcation parameter, and the subscript $t$ denotes partial differentiation with respect to that variable.
@@
Figure 3: Snakes and ladders bifurcation diagrams are a familiar feature of localized patterns in one spatial dimension. The diagram here is for the 1D SHE (1.2) with $\nu=1.6$ and features two intertwined branches of localized solutions: one with a maximum in the middle (blue) and one with a minimum (red). These branches are connected by the asymmetric ‘ladder’ states (green). As one ascends the diagram, the rolling localized pattern grows by adding new rolls on both sides, as exemplified with the profiles taken at successive saddle-node bifurcations on the curve.
In one spatial dimension ($d=1$) the SHE supports stable “roll” patterns for $0<\mu\ll 1$ and $\nu>\sqrt{27/38}$ that resemble sinusoidal functions. Since both rolls and the trivial state $U=0$ are stable in the same parameter regime, an interesting experiment one might undertake is to consider an initial condition in the SHE that is the roll pattern on a compact set and the trivial state on the rest. Now, if the energy of a roll pattern is larger or smaller than 00, the energy of the trivial state $U=0$, then one might intuit that the region of localization (the rolls) contracts or expands as time goes on. For a fixed $\nu$, the energy of a roll pattern will vary with $\mu$ leaving, in general, only a single point of perfect matching where it also has zero energy. This point is referred to as the Maxwell point and it occurs at $\mu=0.2004$ when $\nu=1.6$. This argument seems to indicate that localized roll patterns should only be expected at the Maxwell point, while the numerics presented Figure 3 indicate otherwise. The reason we find such a discrepancy is that the above intuition does not account for the transition between $U=0$ and the localized roll components. The energy stored at these interfaces allows one to observe localized roll patterns in an open parameter interval about the Maxwell point. Pomeau [196] was the first to postulate qualitatively that the pattern could ‘pin’ the interfaces and create an open region in parameter space for stationary patterns to co-exist.
@@
What is not covered? As with all reviews, we have had to make some decisions on what to review and not. The literature on localized patterns is vast, covering decades worth of research. In terms of applications, we only give an overview of some of the main areas, instead focusing on the mathematical aspects. In terms of mathematical analysis and techniques, we do not cover localized and domain-covering patterns occurring near the singular limit (though in §2.4.5 we do describe how the localized patterns change their behavior as they move closer to the singular limit), invasion pattern-forming fronts (though we do link to this literature in §3.1), numerical methods, nonlocal models, and patterns forming in the wake of an invasion front. Our focus is on localized patterns, and so we direct the reader to Cross and Hohenberg’s 1993 review [73] for a wider range of patterns than just those considered here. Furthermore, we only review mathematical concepts at the surface level, while more in-depth mathematical treatments can be found in books by Hoyle [121] and Schneider & Uecker [210].
@@
Two of the most fundamental concepts of studying pattern formation are finite-wavenumber instabilities and the Swift–Hohenberg equation. Interestingly, both of these concepts can be traced back to Alan Turing and his work on morphogenesis. In honor of Turing’s contribution we typically refer to any finite-wavenumber instability in a system of RD equations as a Turing instability , but as a historical note similar investigations were undertaken well before Turing in the fluids literature by Rayleigh [199] and Taylor [221], among others.
@@ 2.1.1 Turing’s Contribution
where $D_{u}$ and $D_{v}$ are the diffusion coefficients of $u$ and $v$, respectively, and subscripts on the variables $u$ and $v$ denote partial differentiation with respect to time $t\geq 0$ and space $x\in\mathbb{R}$.
@@ 2.1.2 General Pattern-Forming Bifurcations
where $\sigma,\nu$ are real parameters, and arose from work in his last years developing a mathematical theory for the formation of patterns in daisies [79]. Similarly, Swift and Hohenberg’s original derivation of the equation that now bears their name had a nonlocal term that has been all but forgotten in the literature [219]. For the rest of this section (and the paper in general), we will concentrate on variants of the SHE that do not include nonlocal terms since this is where much of the mathematical progress has been made. A brief exception will be a quick review of localized patterns in nonlocal equations found in §2.4.2. Furthermore, the local versions of the SHE have a lot in common with RD systems, and so much much of the discussion below also applies to equations of the form (2.1). We discuss extensions of the theory to the general RD setting in §2.4.
@@ 2.3.1 Weakly Nonlinear Analysis
Going to higher-order in the Ginzburg–Landau amplitude equation expansion is still unable to capture the snaking behavior. An alternative method is to use exponential asymptotics. This approach focuses on the region near the co-dimension 2 point where the Turing bifurcation turns from being sub- to supercritical, at which point the width of the snaking is exponentially small and the snaking patterns are small in amplitude. In this case one attempts to explain the snaking bifurcations by expanding around the Maxwell point of the front solutions found in (2.12). Kozyreff and Chapman [147, 65] carried out exponential asymptotics for the quadratic-cubic Swift-Hohenberg equation of the form
@@ 2.3.2 Snakes, Ladders, and Isolas of Localized Patterns
We begin by recalling the recast first-order ODE system (2.10), coming from the original fourth-order time-independent SHE ODE (2.8). Recall that (2.10) relates the existence of steady-state solutions to the SHE as orbits of (2.10) that remain bounded for all $x\in\mathbb{R}$, while localization requiring $U(x)\to 0$ as $x\to\pm\infty$ in the fourth-order ODE (2.8) manifests itself as the condition $(u_{1}(x),u_{2}(x),u_{3}(x),u_{4}(x))\to(0,0,0,0)$ as $x\to\pm\infty$ in (2.10). Thus, we see how the terminology homoclinic snaking arises, as we have reduced ourselves to searching for homoclinic orbits to the spatial ODE (2.10). Moreover, the ‘rolling’ portion of localized states to the SHE can be interpreted as the desired homoclinic orbit undergoing an intermediary winding around a periodic orbit before mounting a return back to a neighborhood of the trivial equilibrium state. Such periodic orbits are exactly the roll patterns discussed earlier in this section since spatial periodicity in (2.8) is equivalent to a periodic solution of (2.10).
Due to the pattern-forming Turing bifurcation at $\mu=0$, for certain fixed $(\mu,\nu)$ the SHE admits a one-parameter family of spatially-periodic, steady-state ‘roll’ patterns that can be parameterized by their period. Intuitively, localized roll patterns as illustrated in Figure 3 can be obtained by truncating one of these periodic profiles over a finite patch $x\in[-L,L]$ and then gluing on the homogeneous rest state $U=0$ for $|x|>L$. Identifying exactly which roll pattern should be used for this gluing procedure can be done by noticing that (2.8) admits the conserved quantity
@@
The above quantity is conserved pointwise in $x$ for solutions $U(x)$ to (2.8) and so any roll pattern that is matched with $U=0$ to form a localized state must lie in $\mathcal{H}(U,\mu,\nu)=\mathcal{H}(0,\mu,\nu)=0$. As there is typically only one roll pattern in this zero-level set, this leads to a selection principle for localized roll patterns away from their onset at $\mu=0$ which was discussed in the previous subsection.
@@
Figure 9: The work of Beck et al. [30] uses spatial dynamics to frame localized roll patterns as homoclinic orbits in the ODE (2.10). Left: A front pattern (bottom) is used to transit between the background state $U=0$ and the roll pattern (a periodic orbit) in phase-space (top). Middle: Ascending the snaking bifurcation diagram is equivalent to extending the length of the localization plateau. Right: Localized patterns are homoclinic to $U=0$ and wrap around the periodic orbit, with more wraps leading to a longer plateau.
We now turn our attention back to the case when localized roll patterns undergo homoclinic snaking. This spatial dynamics approach to obtain the existence of the desired homoclinic orbits to (2.10) requires assuming the existence of a heteroclinic connection between a periodic orbit and $\mathbf{u}=0$ [30, 8, 141, 207]. In practice this assumption can be verified numerically. Such heteroclinic connections are pinned front solutions to (2.8) and reversible symmetry implies that a connection from a roll to $U=0$ exists if and only if a connection from $U=0$ to the roll exists. In the geometry of the 4-dimensional phase-space of (2.10), these heteroclinic orbits provide a method of transiting in phase-space from $\mathbf{u}=0$ to the periodic orbit and back, as illustrated in Figure 9. Beck et al. [30] make the hypothesis that the parameter values $\mu$ at which a heteroclinic connection exists in (2.10) can be parameterized by the phase along the periodic orbit, $\varphi$, where the connection enters a tubular neighborhood of the periodic orbit through $\mu=z(\varphi)$. The result is that symmetric homoclinic orbits that spend time $2L$ in the neighborhood of the periodic orbit exist at a parameter value $\mu$ if and only if
@@
Figure 10: Solutions in $(L,\varphi)$-space satisfying (2.21) can be found graphically by identifying two intersection points with $z$ along a fixed horizontal line whose abscissas give $L\pm\varphi$. Moving the horizontal line up and down generates the entire bifurcation branch, as depicted in the center panel with the resulting bifurcation curve on the right. If $\varphi$ reaches $\pi$ the branch terminates on a second snaking curve, as in Figure 3.
@@
Extensions of the work in [30] abound. For example, [8] shows that if the Floquet exponents of the periodic orbits are negative then all localized states must lie along isolas, regardless of the front/back bifurcation structure. Burke and Dawes examined localized states in a non-variational SHE model and compared them to the SHE studied herein [53]. Various symmetry-breaking results have been documented as well [142, 165, 207], while PDE stability of localized solutions was examined numerically in [54] and analytically in [166]. Knobloch, Uecker, and Wetzel further examined a variant of the SHE with a degree 7 polynomial nonlinearity to find the existence of localized solutions that replace the background state $U=0$ with another smaller amplitude roll pattern [140]. Using the above spatial dynamical framework one can interpret these defect patterns as trajectories that are homoclinic to a periodic orbit but spend an intermediary portion of the orbit wrapping around a different one. All of the above spatial dynamics analysis is therefore applicable to understand these defect patterns as well.
@@ 2.4.1 General Reaction-Diffusion Systems
The spatial dynamical analysis of Beck et al. [30, §6.2] and Knobloch and Wagenknecht [143] can be carried over to higher dimensional ODEs with and without Hamiltonian structure. One major exception is that the loss of a Hamiltonian or conservative structure in the resulting spatial dynamical system leads the asymmetric ladder states to become dynamic. That is, they move with nonzero speed either to the left or right in the spatial domain while maintaining their localized profile.

### v2

# Localized Patterns

@@ 1 Introduction
In this review, we will focus on structures that are localized in one or more spatial dimensions embedded in a quiescent state and possess some non-trivial fine-scale regular structure, which we call a pattern. Figure 1 provides examples of the types of localized patterns we will concentrate on that occur in both continuum and discrete systems. Throughout this review, we will see that the field has been developed with numerical, experimental, and analytical techniques that work in tandem to better understand localization mechanisms. We highlight several reviews that came before this one [76, 143, 142, 66, 172], but note their focus is more rooted in the physics literature or primarily restricted to localized patterns in one spatial dimension. What differentiates this review from those that came before is that we seek to highlight the mathematical development of the subject. This includes how specific equations have been investigated for localized patterns and providing overviews of the myriad of clever mathematical and computational techniques that have been developed to study these multi-dimensional patterns.
Localized patterns can be found in a wide range of application areas from fluids, quantum theory, biology, mechanics, sociology, to nonlinear optics. They come in the form of an oasis of vegetation in an arid landscape, a hotspot of criminal activity, or a laser pulse in plasmas. Because they arise in so many different applications, they are often referred to by different names which have historical relevance to the field. This includes solitary waves in fluid dynamics, convectons in binary fluid convection, and fairy circles in vegetation models, to name a few. To ease the reader’s journey into the literature on localized patterns, we will provide references and nomenclature for each specific type of pattern as they occur in this review. This will help to connect the specific pattern to specific applications. However, we note that while the study of these patterns is motivated by application, our goal is to review all things mathematical in their investigation. This means primarily relying on phenomenological equations that describe macroscopic behavior and attempting to understand the fundamental mechanisms that drive their formation.
@@
where $\mathbf{u}=\mathbf{u}(\mathbf{x},t)\in\mathbb{R}^{n}$, $\mathbf{D}$ is an $n\times n$ diagonal ‘diffusion’ matrix whose entries are all positive, $\Delta$ is the Laplace operator on $\mathbb{R}^{d}$, $\mathbf{f}$ is a smooth nonlinearity, $\mu$ is a bifurcation parameter, and the subscript $t$ denotes partial differentiation with respect to that variable.
@@
Figure 3: Snakes and ladders bifurcation diagrams are a familiar feature of localized patterns in one spatial dimension. The diagram here is for the 1D SHE (1.2) with $\nu=1.6$ and features two intertwined branches of localized solutions: one with a maximum in the middle (blue) and one with a minimum (red). These branches are connected by the asymmetric ‘ladder’ states (green). As one ascends the diagram, the localized pattern grows by adding new spatially periodic patterns (SPPs) on both sides, as exemplified with the profiles taken at successive saddle-node bifurcations on the curve.
In one spatial dimension ($d=1$) the SHE supports stable spatially periodic patterns (SPPs) for $0<\mu\ll 1$ and $\nu>\sqrt{27/38}$ that resemble sinusoidal functions. Since both the SPP and the trivial state $U=0$ are stable in the same parameter regime, an interesting experiment one might undertake is to consider an initial condition in the SHE that is the SPP on a compact set and the trivial state on the rest. Now, if the energy of a SPP is larger or smaller than 00, the energy of the trivial state $U=0$, then one might intuit that the region of localization (the SPP) contracts or expands as time goes on. For a fixed $\nu$, the energy of a SPP will vary with $\mu$ leaving, in general, only a single point of perfect matching where it also has zero energy. This point is referred to as the Maxwell point and it occurs at $\mu=0.2004$ when $\nu=1.6$. This argument seems to indicate that localized SPPs should only be expected at the Maxwell point, while the numerics presented Figure 3 indicate otherwise. The reason we find such a discrepancy is that the above intuition does not account for the transition between $U=0$ and the localized SPP components. The energy stored at these interfaces allows one to observe localized SPPs in an open parameter interval about the Maxwell point. Pomeau [203] was the first to postulate qualitatively that the pattern could ‘pin’ the interfaces and create an open region in parameter space for stationary patterns to co-exist.
@@
What is not covered? As with all reviews, we have had to make some decisions on what to review and not. The literature on localized patterns is vast, covering decades worth of research. In terms of applications, we only give an overview of some of the main areas, instead focusing on the mathematical aspects. The focus of the review is not on solitons or solitary waves in integrable systems though a lot of the techniques described below can be applied to those systems. Instead, this review will focus on unforced non-equilibrium thermodynamical systems such as the RD systems and SHE. Additionally, we focus on the existence of localized patterns and not their stability; the literature surrounding the (in)stability of localized patterns is similarly extensive and worthy of its own review. In terms of mathematical analysis and techniques, we do not cover localized and domain-covering patterns occurring near the singular limit where one of the diffusion coefficients in the RD system is very small (though in §2.4.5 we do describe how the localized patterns change their behavior as they move closer to the singular limit), invasion pattern-forming fronts (though we do link to this literature in §3.1), numerical methods, nonlocal models, and patterns forming in the wake of an invasion front. Our focus is on localized patterns, and so we direct the reader to Cross and Hohenberg’s 1993 review [75] for a wider range of patterns than just those considered here. Furthermore, we only review mathematical concepts at the surface level, while more in-depth mathematical treatments can be found in books by Hoyle [126] and Schneider & Uecker [217].
@@
Two of the most fundamental concepts of studying pattern formation are finite-wavenumber instabilities and the Swift–Hohenberg equation. Interestingly, both of these concepts can be traced back to Alan Turing and his work on morphogenesis. In honor of Turing’s contribution we typically refer to any finite-wavenumber instability in a system of RD equations as a Turing instability , but as a historical note similar investigations were undertaken well before Turing in the fluids’ literature by Rayleigh [206] and Taylor [228], among others.
@@ 2.1.1 Turing’s Contribution
where $D_{u},D_{v}>0$ are the diffusion coefficients of $u$ and $v$, respectively, and subscripts on the variables $u$ and $v$ denote partial differentiation with respect to time $t\geq 0$ and space $x\in\mathbb{R}$.
@@ 2.1.2 General Pattern-Forming Bifurcations
where $\sigma,\nu$ are real parameters, and arose from work in his last years developing a mathematical theory for the formation of patterns in daisies [81]. Similarly, Swift and Hohenberg’s original derivation of the equation that now bears their name had a nonlocal term that has been all but forgotten in the literature [226]. For the rest of this section (and the paper in general), we will concentrate on variants of the SHE that do not include nonlocal terms since this is where much of the mathematical progress has been made. A brief exception will be a quick review of localized patterns in nonlocal equations found in §2.4.2. Furthermore, the local versions of the SHE have a lot in common with RD systems, and so much of the discussion below also applies to equations of the form (2.1). We discuss extensions of the theory to the general RD setting in §2.4.
@@ 2.3.1 Weakly Nonlinear Analysis
Going to higher-order in the Ginzburg–Landau amplitude equation expansion is still unable to capture the snaking behavior. An alternative method is to use exponential asymptotics [152, 67, 83, 82]. This approach focuses on the region near the co-dimension 2 point where the Turing bifurcation turns from being sub- to supercritical, at which point the width of the snaking is exponentially small and the snaking patterns are small in amplitude. In this case one attempts to explain the snaking bifurcations by expanding around the Maxwell point of the front solutions found in (2.12), and so the resulting analysis is highly dependent on the choice of model. Kozyreff and Chapman [152, 67] carried out exponential asymptotics for the quadratic-cubic Swift-Hohenberg equation of the form
@@ 2.3.2 Snakes, Ladders, and Isolas of Localized Patterns
The weakly nonlinear analysis in the previous section is highly model dependent and only covers small amplitude patterns. In this section we look at the general geometric arguments for homoclinic snaking that applies to large amplitude localized patterns. We begin by recalling the recast first-order ODE system (2.10), coming from the original fourth-order time-independent SHE ODE (2.8). Recall that (2.10) relates the existence of steady-state solutions to the SHE as orbits of (2.10) that remain bounded for all $x\in\mathbb{R}$, while localization requiring $U(x)\to 0$ as $x\to\pm\infty$ in the fourth-order ODE (2.8) manifests itself as the condition $(u_{1}(x),u_{2}(x),u_{3}(x),u_{4}(x))\to(0,0,0,0)$ as $x\to\pm\infty$ in (2.10). Thus, we see how the terminology homoclinic snaking arises, as we have reduced ourselves to searching for homoclinic orbits to the spatial ODE (2.10). Moreover, the SPP portion of localized states to the SHE can be interpreted as the desired homoclinic orbit undergoing an intermediary winding around a periodic orbit before mounting a return back to a neighborhood of the trivial equilibrium state. Such periodic orbits are exactly the SPPs discussed earlier in this section since spatial periodicity in (2.8) is equivalent to a periodic solution of (2.10).
Due to the pattern-forming Turing bifurcation at $\mu=0$, for certain fixed $(\mu,\nu)$ the SHE admits a one-parameter family of spatially-periodic, steady-state SPPs that can be parameterized by their period. Intuitively, localized SPPs as illustrated in Figure 3 can be obtained by truncating one of these periodic profiles over a finite patch $x\in[-L,L]$ and then gluing on the homogeneous rest state $U=0$ for $|x|>L$. Identifying exactly which SPP should be used for this gluing procedure can be done by noticing that (2.8) admits the conserved quantity
@@
The above quantity is conserved pointwise in $x$ for solutions $U(x)$ to (2.8) and so any SPP that is matched with $U=0$ to form a localized state must lie in $\mathcal{H}(U,\mu,\nu)=\mathcal{H}(0,\mu,\nu)=0$. As there is typically only one SPP in this zero-level set, this leads to a selection principle for localized SPPs away from their onset at $\mu=0$ which was discussed in the previous subsection.
@@
Figure 9: The work of Beck et al. [30] uses spatial dynamics to frame localized spatially periodic patterns (SPPs) as homoclinic orbits in the ODE (2.10). Left: A front pattern (bottom) is used to transit between the background state $U=0$ and the SPP in phase-space (top). Middle: Ascending the snaking bifurcation diagram is equivalent to extending the length of the localization plateau. Right: Localized patterns are homoclinic to $U=0$ and wrap around the periodic orbit, with more wraps leading to a longer plateau.
We now turn our attention back to the case when localized SPPs undergo homoclinic snaking. This spatial dynamics approach to obtain the existence of the desired homoclinic orbits to (2.10) requires assuming the existence of a heteroclinic connection between a periodic orbit and $\mathbf{u}=0$ [30, 8, 146, 214]. In practice this assumption can be verified numerically. Such heteroclinic connections are pinned front solutions to (2.8) and reversible symmetry implies that a connection from a SPP to $U=0$ exists if and only if a connection from $U=0$ to the SPP exists. In the geometry of the 4-dimensional phase-space of (2.10), these heteroclinic orbits provide a method of transiting in phase-space from $\mathbf{u}=0$ to the periodic orbit and back, as illustrated in Figure 9. Beck et al. [30] make the hypothesis that the parameter values $\mu$ at which a heteroclinic connection exists in (2.10) can be parameterized by the phase along the periodic orbit, $\varphi$, where the connection enters a tubular neighborhood of the periodic orbit through $\mu=z(\varphi)$. The result is that symmetric homoclinic orbits that spend time $2L$ in the neighborhood of the periodic orbit exist at a parameter value $\mu$ if and only if
@@
Figure 10: Solutions in $(L,\varphi)$-space satisfying (2.21) can be found graphically by identifying two intersection points with $z$ along a fixed horizontal line whose abscissas give $L\pm\varphi$. Moving the horizontal line up and down generates the entire bifurcation branch, as depicted in the center panel, with the resulting bifurcation curve on the right. If $\varphi$ reaches $\pi$ the branch terminates on a second snaking curve, as in Figure 3.
@@
Extensions of the work in [30] abound. For example, [8] shows that if the Floquet exponents of the periodic orbits are negative then all localized states must lie along isolas, regardless of the front/back bifurcation structure. Burke and Dawes examined localized states in a non-variational SHE model and compared them to the SHE studied herein [54]. Various symmetry-breaking results have been documented as well [147, 170, 214], while PDE stability of localized solutions was examined numerically in [55] and analytically in [171]. Knobloch, Uecker, and Wetzel further examined a variant of the SHE with a degree 7 polynomial nonlinearity to find the existence of localized solutions that replace the background state $U=0$ with another smaller amplitude SPP [145]. Using the above spatial dynamical framework one can interpret these defect patterns as trajectories that are homoclinic to a periodic orbit but spend an intermediary portion of the orbit wrapping around a different one. All of the above spatial dynamics analysis is therefore applicable to understand these defect patterns as well.
@@ 2.4.1 General Reaction-Diffusion Systems
The spatial dynamical analysis of Beck et al. [30, §6.2] and Knobloch and Wagenknecht [148] can be carried over to higher dimensional ODEs with and without Hamiltonian structure. One major exception is that the loss of a Hamiltonian or conservative structure in the resulting spatial dynamical system leads the asymmetric ladder states to become dynamic. That is, they move with nonzero speed either to the left or right in the spatial domain while maintaining their localized profile [125].